Given the linear equation y= 6x+2 find the y-coordinates of the points(-1),(-2),& (1). Please show all work. Plot these points & gaph the linear equation.
Answer by mananth(16949)

y= 6x+2 find the y-coordinates of the points(-1),(-2),& (1).
-1
y=-6+2=-4 (-1,-4)
-2
y=-12+2=-10 (-2,-10)
1
y=6+2=8 (1,8)